OpenStudy (anonymous):

For homework, Fred has to read 30 minutes daily. It takes him 2 minutes per page, and he spends 4 minutes viewing graphics in the book. If x represents the number of pages he reads, then 2x + 4 ≥ 30 represents the time he reads each day. What is the fewest number of pages Fred must read to complete his homework?

OpenStudy (anonymous):

i believe its 13

OpenStudy (anonymous):

yea i thot so

OpenStudy (danjs):

they gave you the inequality here, 2 min/page*(#pages) + 4 min >= 30 minutes total

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i have another

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

13

OpenStudy (anonymous):

Which equation best represents the situation? Cara served pizza and punch at a party. Each pizza fed 6 people, and each guest received a glass of punch. If a glass of punch cost 50¢, and Cara spent $2.50 for each guest, what was the cost of each pizza?

OpenStudy (danjs):

solve for the range of x, the pages

OpenStudy (anonymous):

ur too late dan lol

OpenStudy (danjs):

well they don't give you the inequality this time,,

OpenStudy (anonymous):

Which equation best represents the situation? Cara served pizza and punch at a party. Each pizza fed 6 people, and each guest received a glass of punch. If a glass of punch cost 50¢, and Cara spent $2.50 for each guest, what was the cost of each pizza?

OpenStudy (danjs):

tso the concept has to be put together

OpenStudy (anonymous):

can someone help me with this problem

OpenStudy (anonymous):

Which equation best represents the situation? Cara served pizza and punch at a party. Each pizza fed 6 people, and each guest received a glass of punch. If a glass of punch cost 50¢, and Cara spent $2.50 for each guest, what was the cost of each pizza?

OpenStudy (anonymous):

it says i gotta log in... the best way i can help you is by saying she only bought one pizza because she only had 5 guests

OpenStudy (danjs):

Each guest cost $2.50 total, that is the same cost as 1/6 of a pizza, and one glass of drink

OpenStudy (danjs):

put that into an equation 2.5 = p/6 + 0.5 balancing the $$

OpenStudy (anonymous):

oh then the pizza would be twelve dollars?

OpenStudy (anonymous):

a.x/6+.50=2.50 b.6x+.50x=2.50 c.x/6+50=2.50 d.6x-2.50=.50

OpenStudy (danjs):

1 guest cost 2.50 and consumed 1/6 a pizza and one drink 2.50 = x/6 + 0.5

OpenStudy (anonymous):

so it would be a

OpenStudy (danjs):

yes,

OpenStudy (anonymous):

oh yeah lol..

OpenStudy (anonymous):

okay thx i have a couple more if you guys can help?

OpenStudy (danjs):

The main point is to read the thing a couple times, summarize it, and convert that to equation(s)

OpenStudy (anonymous):

suuure

OpenStudy (danjs):

If you can keep track of units well, that makes it much easier

OpenStudy (anonymous):

Which inequality best represents the situation? A pine tree grows at a rate of 1.5 feet per year. James plants a 2-foot tall pine tree in his yard. How many years will it take for the tree to be at least 14 feet tall?

OpenStudy (anonymous):

okay dan

OpenStudy (anonymous):

A. 2x + 1.5 ≥ 14 B. 2x – 1.5 ≤ 14 C. 1.5x + 2 ≥ 14 D. 1.5x – 2 ≤ 14

OpenStudy (anonymous):

i think d

OpenStudy (danjs):

Here is my thought, 1)the tree grows at a constant rate, 1.5 ft / 1 yr, it must be a linear relationship

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so we divide

OpenStudy (danjs):

for each increase in yrs, the height grows the same amount, constant slope for a line

OpenStudy (anonymous):

i think its c

OpenStudy (danjs):

says it starts at 2 ft tall, that is the value at time of zero, and the y intercept for the line

OpenStudy (anonymous):

right so that means that we need to dive those 2 numbers

OpenStudy (anonymous):

okay

OpenStudy (danjs):

they give you 1 point, and the slope, you can make an equation for a line

OpenStudy (anonymous):

its c

OpenStudy (danjs):

y = m*x+b y = 1.5x + 2

OpenStudy (anonymous):

hwo is that mega

OpenStudy (anonymous):

oh so it is c y

OpenStudy (anonymous):

exactly dans equation y = 1.5x + 2

OpenStudy (anonymous):

ohhhhh right

OpenStudy (anonymous):

okay

OpenStudy (danjs):

now that you know the equation for the line, y = 1.5*x + 2 they want to know when it will be at least 30 ft tall? the 1.5x + 2 has to be greater than or equal to 30 ft

OpenStudy (anonymous):

is that it?

OpenStudy (danjs):

yes, that is how i would think thorugh it, basically

OpenStudy (anonymous):

You want to find out how many cards were in Mona's deck at the beginning. Which situation about her game best represents the equation x/4+2=10 ?

OpenStudy (anonymous):

Mona gave 4 players an equal number of game cards and then put 2 cards back in the box. After that, the first player had 10 cards. B. Mona gave 4 players an equal number of cards and then gave the first player 2 extra cards. After that, the first player had 10 cards. C. Mona gave 4 players an equal number of cards and then gave each player 2 additional cards. After that, each player had 10 cards. D. Mona gave 4 players an equal number of cards and then gave 1 more card to 2 players. After that, the first player had 10 cards.

OpenStudy (anonymous):

I think b

OpenStudy (anonymous):

no c

OpenStudy (anonymous):

right?

OpenStudy (danjs):

from the question it is hard to tell, since there are no units for x and 10, cards vs players

OpenStudy (danjs):

so you have to read the responses, and figure that

OpenStudy (anonymous):

hmmmm?

OpenStudy (anonymous):

could u help me with that?

OpenStudy (danjs):

x is the number of cards in the full deck

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so we would subtract

OpenStudy (anonymous):

without x

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

come one mathlete

OpenStudy (danjs):

sorry, had to step away a second...

OpenStudy (anonymous):

thats okay

OpenStudy (danjs):

the 4 is the number of players that was given, ...

OpenStudy (anonymous):

right

OpenStudy (danjs):

the total cards x, was divided by the number of players, x/4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so substract?

OpenStudy (danjs):

thif you look at the units there, you get cards per player cards / player the units have to be the same on both sides of the equation, so that constant 2, and 10 will also be cards/player

OpenStudy (danjs):

cards / player + cards / player = cards / player

OpenStudy (anonymous):

so the answer is c right? or

OpenStudy (danjs):

sorry, i will restart, was distracted, and this is wrong.

OpenStudy (anonymous):

what is ?

OpenStudy (anonymous):

is is a ? do u know the answer?

OpenStudy (anonymous):

talk to meh

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

hey umm? so i kinda dont have much time till i have to go so if you could just ya know help meh that would be awesome lol

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

right?